OpenStudy (anonymous):

Pam scored 78 on a test that had 4 fill-in questions worth 7 points each and 24 multiple-choice questions worth 3 points each. She had one fill-in question wrong. How many multiple-choice questions did Pam get right?

OpenStudy (anonymous):

Let's start with 100, a perfect score. She had to lose 22 points in order to get her 78 grade. We need to account for all 22 of these points. So: -We know for a fact she got one fill-in question wrong, which subtracts 7 points. This means 15 more points (22-7) need to be accounted for. -How many multiple choice questions will add up to 15 points? 15/3 = 5. She got 5 multiple choice questions wrong, which would mean that she got 22 points total off, which would mean she got a 78%.

OpenStudy (campbell_st):

100 = 4*7 + 24*3 but im your question 1 7 mark was wrong let x be the number of correct multiple choice questions so 78 = 3*7 + 3*x 78 = 21 + 3x subtract 21 from both sides 57 = 3x divide both sides by 3 19 = x Pam correctly answered 19 multiple choice questions.